OpenStudy (anonymous):

lim x approaches infinity of ln (e^x -1) - x

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} e ^{x} - 1 - x\]

OpenStudy (anonymous):

is that the ques?

OpenStudy (experimentx):

isn't this obvious?? when x grows every large ... e^x completely dominates -1

OpenStudy (anonymous):

no this is not the question, it is the limit of; ln(e^x +1)-x

OpenStudy (experimentx):

what ever there is e^x+1 =e^x as x goes very large.

OpenStudy (experimentx):

the answer is 0 ... just by looking!!

OpenStudy (experimentx):